A dairy keeps two kinds of milk on hand, skim milk has 0.2% butterfat and whole milk has 3.4% butterfat. How many gallons of each type of milk does a company need to produce 220 gallons of 1% milk for a gorcery store?
Answer by ankor@dixie-net.com(22740)

A dairy keeps two kinds of milk on hand, skim milk has 0.2% butterfat and whole milk has 3.4% butterfat.
How many gallons of each type of milk does a company need to produce 220 gallons of 1% milk for a grocery store?
:
let x = amt of 3.4% milk
the total is to be 220 gal, therefore
(220-x) = amt of .2% milk
:
The amt of butterfat equation
.034x + .002(220-x) = .01(220)
.034x + .44 - .002x = 2.20
.034x - .002x = 2.20 - .44
.032x = 1.76
x = 1.76/.032
x = 55 Gal of 3.4% mile required
then
220 - 55 = 165 gal of .2% required